shift summary for EVE 9/13/99

Richard Ford (rickford@fnal.gov)
Tue, 14 Sep 1999 00:12:43 -0500 (CDT)

Shift Summary for Evening Shift, 09/13/99

On Shift: Julie Whitmore, Rick Ford

Hours of beam available at all intensities: 8.5
Hours of beam available at nominal intensity: 8.5
Hours of E832 data (special runs not included): 6.2
Hours of special runs (laser scans, muon runs...): 0.00

Accounting of lost time:

========================================================================

Runs ended this shift: None

Tapes brought to FCC this shift: None


========================================================================

Accesses during shift: Access to fix reg. HV vacuum interlock

========================================================================

Overview of shift: 832 data taking at 4.0 E12
Reg. HV probs at the beginning with smoooth sailing
thereafter

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15:30 Inherit shift with daq down due to A/C problems and
missing vacuum interlock permit for reg. HV.

14516 19 17:50 After 2 accesses and consultation w/Dave Pushka at
home it is determined that we cannot fix the vacuum
interlock problem quickly and we will just have to
run without it and watch the vacuum manually until
it can be fixed during the shutdown.
A/C problem was fixed by the maintenance contractor.
Beam returns.

90 19:39 Live_time error; possibly FPED process conflict?

19:40 Leonard calls to say hall is running colder than
normal by about 2 degrees C. He will be here in
the morning to investigate.

150 20:50 Julie notices that the cosmics counters are
running hot due to the pedestals shifting 10-15 cts.
Will leave a note for the next crew to update
constants when the run ends.

244 22:58 DYC in laser crate full, kfrend.

292 24:00 Shift ends.


========================================================================
Follow up Notes on Previous Shift

Fixed all of Doug's disasters
========================================================================
Items to Follow up on Next Shift

Muon run and polarity flip, download ADGN constants

Watch for MA scaler or vacuum alarm - reg HV's need to be turned off
manually. See main alarm screen for instructions.

If humidity alarms go off (humidity above 10% in the blockhouse)
call Leonard Nelson. His home number is 630-231-1416; pager
888-750-0702.